# Excitonic Theory of the Ultrafast Optical Response of 2D-Quantum-Confined Semiconductors at Elevated Densities

### v1

@@ IV Excitonic Rabi Oscillations
Figure 10: Dynamics of the total exciton density $N=\frac{1}{\mathcal{A}}|P{\vphantom{P}}_{1s}^{K,\uparrow}|^{2}+\frac{1}{\mathcal{A}}\sum_{\mathbf{Q}}N\vphantom{N}_{1s,\mathbf{Q}}^{K,\uparrow,K,\uparrow}$ at the $K$ valley for circular excitation (a) and linear excitation (b) in a h-BN-encapsulated MoSe2 monolayer.
In Fig. 10, we depict the dynamics of the total exciton density at the $K$ valley from Eq. (46), i.e., $\xi=\xi^{\prime}=K,\uparrow$, for circular (a), cf. also Fig. 7(j), and linear (b) excitation. For the linear-excitation case, we include coherent transitions at the $K$ and $K^{\prime}$ valleys $P{\vphantom{P}}_{1s}^{K}$, $P{\vphantom{P}}_{1s}^{K^{\prime}}$ as well as intravalley incoherent occupations $N\vphantom{N}_{1s,\mathbf{Q}}^{K,K}$, $N\vphantom{N}_{1s,\mathbf{Q}}^{K^{\prime},K^{\prime}}$ and neglect any intervalley occupations $N\vphantom{N}_{1s,\mathbf{Q}}^{K,K^{\prime}}$$N\vphantom{N}_{1s,\mathbf{Q}}^{K^{\prime},K}$. While in circular excitation, some Rabi flops of the coherent exciton density remain, cf. Fig. 8 and the discussion above, in linear excitation, even the Rabi flops in the coherent exciton density are almost totally suppressed. The reason is the larger two-exciton/exciton-two-exciton phase space: In circular excitation, only symmetric (++) intravalley two-exciton continua $B\vphantom{B}_{+,\zeta}^{K,K,K,K}$ and exciton-two-exciton continua $Z\vphantom{Z}_{+,\zeta,\mathbf{Q}}^{K,K,K,K,K,K}$ are induced. In linear excitation, the $K$ valley and the $K^{\prime}$ valley are equally excited, so that additional symmetric (++) and antisymmetric (-) intervalley two-exciton continua $B\vphantom{B}_{\pm,\zeta}^{K,K^{\prime},K^{\prime},K}$$B\vphantom{B}_{\pm,\zeta}^{K^{\prime},K,K,K^{\prime}}$ and exciton-two-exciton continua $Z\vphantom{Z}_{\pm,\zeta,\mathbf{Q}}^{K^{\prime},K^{\prime},K,K,K^{\prime},K^{\prime}}$$Z\vphantom{Z}_{\pm,\zeta,\mathbf{Q}}^{K,K,K^{\prime},K^{\prime},K,K}$ are induced. These yield a dynamic coupling between the excitonic transitions at the $K$ valley $P{\vphantom{P}}_{1s}^{K}$ and at the $K^{\prime}$ valley $P{\vphantom{P}}_{1s}^{K^{\prime}}$, hence further increasing the excitation-induced dephasing. Moreover, the formation of a bound ($b$) biexciton $B\vphantom{B}_{-,\zeta=b}^{K,K^{\prime},K^{\prime},K}$$B\vphantom{B}_{-,\zeta=b}^{K^{\prime},K,K,K^{\prime}}$ and bound exciton-biexciton $Z\vphantom{Z}_{-,\zeta=b,\mathbf{Q}}^{K^{\prime},K^{\prime},K,K,K^{\prime},K^{\prime}}$$Z\vphantom{Z}_{-,\zeta=b,\mathbf{Q}}^{K,K,K^{\prime},K^{\prime},K,K}$ becomes possible, further enhancing the excitation-induced dephasing. Hence, due to the overall larger number of Coulomb-interaction-induced exciton-exciton correlations and the resulting dynamic coupling of the excitonic transitions $P$ between the $K$ and $K^{\prime}$ valley, the excitation-induced dephasing is larger in linear excitation compared to circular excitation causing a suppression of the Rabi oscillations in the coherent exciton density. This theoretical finding is consistent with recent experiments in a MoSe2 monolayer, where a clear Rabi splitting due to an occupation grating, which is insensible to coherent excitation-induced dephasing, has been measured, but no Rabi oscillations could be observed Schäfer et al. (2025).
@@ V Conclusion
In this work, we developed and discussed a theory for the ultrafast dynamics of Wannier excitons valid from the low-density regime up to an elevated-density regime below the Mott transition, which is unleashed on two distinct materials as proponents of different Coulomb-to-light-matter-interaction regimes: A GaAs quantum well with moderate Coulomb interaction and a MoSe2 monolayer with strong Coulomb interaction compared to light-matter interaction. It has been found, that the excitonic approach reproduces the key physics described by the semiconductor Bloch equations well, which have been proven to reliably model Rabi flops in ultrafast experiments in the moderate Coulomb-interaction regime (GaAs QW) Schülzgen et al. (1999); Binder et al. (1999); Cundiff et al. (1994), while showing major deviations from the SBEs in the strong Coulomb-interaction regime (MoSe2 ML) due to exciton-exciton interaction induced by Coulomb-correlated doublets and triplets, in line with recent experiments Schäfer et al. (2025). Here, in circular excitation, Rabi oscillations are considerably less strong and vanish almost completely in linear excitation. Our findings strongly indicate, that an excitonic approach has to be chosen to simulate the non-linear response of confined semiconductors, as long as the excitation conditions are below the Mott transition and correlated electron-hole pairs dominate. On the other hand, if the excitation conditions are above the Mott transition, where uncorrelated electrons and holes dominate, the expansion in correlated electron-hole pairs breaks down and the semiconductor Bloch equations need to be applied.
@@ Acknowledgements.
Funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – Project No. 420760124 (H.M., A.K.); 556436549 (A.K.). H.M. and A.K. thank Felix Schäfer, Markus Stein and Sangam Chatterjee (Justus-Liebig-Universität Gießen) for fruitful discussions and their measurements, which strongly motivated the theory developed here.

### v2

@@ IV Excitonic Rabi Oscillations
Figure 11: Dynamics of the total exciton density $N=\frac{1}{\mathcal{A}}|P{\vphantom{P}}_{1s}^{K,\uparrow}|^{2}+\frac{1}{\mathcal{A}}\sum_{\mathbf{Q}}N\vphantom{N}_{1s,\mathbf{Q}}^{K,\uparrow,K,\uparrow}$ at the $K$ valley for circular excitation (a) and linear excitation (b) in a h-BN-encapsulated MoSe2 ML.
@@ V Conclusion
In this work, we developed and discussed a theory for the ultrafast dynamics of Wannier excitons valid from the low-density regime up to an elevated-density regime below the Mott transition, which is unleashed on two distinct materials as proponents of different Coulomb-to-light-matter-interaction regimes: A GaAs QW with moderate Coulomb interaction and a MoSe2 ML with strong Coulomb interaction compared to light-matter interaction.
It has been found, that the excitonic approach reproduces the key physics described by the semiconductor Bloch equations well, which have been proven to reliably model Rabi flops in ultrafast experiments in the moderate Coulomb-interaction regime (GaAs QW) [107, 19, 29], while showing major deviations from the SBEs in the strong Coulomb-interaction regime (MoSe2 ML) due to exciton-exciton interaction induced by Coulomb-correlated doublets and triplets, in line with recent experiments [103]. Here, in circular excitation, Rabi oscillations are considerably less strong and vanish almost completely in linear excitation. Our findings strongly indicate, that an excitonic approach has to be chosen to simulate the non-linear response of confined semiconductors, as long as the excitation conditions are below the Mott transition and correlated electron-hole pairs dominate. On the other hand, if the excitation conditions are above the Mott transition [38, 131, 82, 114, 41], where uncorrelated electrons and holes dominate, the expansion in correlated electron-hole pairs breaks down and the semiconductor Bloch equations need to be applied.
@@ Acknowledgements.
Funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – Project No. 420760124 (H.M., A.K.); 556436549 (A.K.). H.M. acknowledges funding by project 21209528 (“proof of trust”). H.M. and A.K. thank Felix Schäfer, Markus Stein and Sangam Chatterjee (Justus-Liebig-Universität Gießen) for fruitful discussions and their measurements, which strongly motivated the theory developed here.
Data Availability Statement
The data supporting the findings of this work are available from the authors within reasonable request.